# Truck Parking Usage Prediction with Decomposed Graph Neural Networks

### v1

@@ Abstract
Truck parking on freight corridors faces various challenges, such as insufficient parking spaces and compliance with Hour-of-Service (HOS) regulations. These constraints often result in unauthorized parking practices, causing safety concerns. To enhance the safety of freight operations, providing accurate parking usage prediction proves to be a cost-effective solution. Despite the existing research demonstrating satisfactory accuracy for predicting individual truck parking site usage, few approaches have been proposed for predicting usage with spatial dependencies of multiple truck parking sites. We present the Regional Temporal Graph Neural Network (RegT-GCN) as a predictive framework for assessing parking usage across the entire state to provide better truck parking information and mitigate unauthorized parking. The framework leverages the topological structures of truck parking site distributions and historical parking data to predict occupancy rates across a state. To achieve this, we introduce a Regional Decomposition approach, which effectively captures the geographical characteristics. We also introduce the spatial module working efficiently with the temporal module. Evaluation results demonstrate that the proposed model surpasses other baseline models, improving the performance by more than $20\%$ compared with the original model. The proposed model allows truck parking sites’ percipience of the topological structures and provides higher performance.
@@ I Introduction
The Federal Motor Carrier Safety Administration (FMCSA) has implemented regulations to decrease fatigue-related safety hazards for truck drivers by restricting the consecutive and total daily driving hours permissible for truck operators. Despite these efforts, truck drivers still spend considerable time searching for safe parking locations beyond their limited legal working hours [1]. This phenomenon highlights the persistent issue of inadequate truck parking facilities, contributing to unauthorized parking behaviors and breaches of parking regulations [2].
@@
The literature identifies increasing truck drivers’ accessibility to real-time parking usage information [3] and drivers’ perceptual interpretations [4, 5] as major factors influencing these parking challenges. According to the survey studies in the region [6], members of the Mid-America Association of State Transportation Officials (MAASTO) from eight states, Indiana, Iowa, Kansas, Kentucky, Michigan, Minnesota, Ohio, and Wisconsin, have collaborated to construct a real-time, multi-state Truck Parking Information Management System (TPIMS) to feed more practical truck parking information for truck drivers [7, 8]. TPIMS, fully operating since January 4th, 2019, provides real-time parking usage information to truck drivers through dynamic message signs, smartphone applications, traveler information websites (e.g., 511 traveler information), and other forms.
@@
In addition to real-time parking site availability information, future usage prediction can better support truck drivers in their decision-making process and route planning. Current prediction models exhibit commendable accuracy in predicting future usage for the subsequent hour [9, 10]. However, it is crucial to recognize that the domain of interest encompasses multiple truck parking sites situated along the highway corridor. Boris and Brewster [1] reported that the most critical aspects of the target parking sites are the proximity to the route/destination followed by the amenity availability. Thus, the primary objective of our prediction models is to capitalize on the wealth of data available from various truck parking sites and leverage their topological structures.
@@
Since 2019, the research team has developed the TPIMS data archiving system (TPIMS-DAS), designed to store effectively and track truck parking information [11]. Leveraging the substantial benefits derived from an abundance of comprehensive and detailed information on truck parking sites, users can extract and aggregate truck parking data from eight states in MAASTO using TPIMS-DAS. It also leads to providing a sufficient number of truck parking data to train reliable models. Using the advantages of wide-range truck parking data with TPIMS-DAS, we fundamentally analyzed truck parking usage patterns. [12]. In this research, we developed a model grounded in spatial dependencies of highway corridors to comprehend how regional relationships of truck parking site locations can contribute to the future usage prediction task.
@@
This study first addresses the topological structures of highway corridors to assess multiple truck parking site usage prediction. We propose a Regional Temporal Graph Neural Network (RegT-GCN) composed of a Graph Neural Network with temporal capability and novel decomposition modules to acquire regional relationships, separated by state (Figure 1). Utilizing the significant parking site information dataset, we could leverage temporal dependencies of the multi-state-wide truck parking usage to acquire more accurate predicted results. We conducted several experiments to evaluate the importance of Regional Decomposition for Graph Neural Networks.
Our technical contributions can be summarized in the following.
  • A spatio-temporal model to predict the multiple truck parking usage.
@@
  • Extensive quantitative evaluation, which demonstrates that our RegT-GCN significantly outperforms state-of-the-art baseline models.
@@ II Literature Review
Parking prediction methods can be classified into two main categories: urban parking prediction, which pertains to parking prediction within cities, and truck parking prediction, which deals with predicting parking usage for trucks. Typically, urban parking prediction methods offer a shorter prediction horizon, ranging from 5 to 15 minutes ahead. In contrast, truck parking prediction provides predictions for up to 1 hour ahead [10]. Regardless of the distinctions in time frames between these parking prediction tasks, it is noteworthy that both tasks often share fundamental concepts and are devised with a broader perspective encompassing diverse parking scenarios.
@@ II-A Statistical Prediction Model for Truck Parking Usage
Bayraktar et al. [13] developed an intelligent parking-management system for commercial vehicles in Florida, analyzed data, and created a prediction model for rest area occupancy with a Kalman filter. As the Kalman filter prediction-correction mechanism can be utilized with historical data and adapted to online status, their model achieved better performance. Sadek et al. [9] identified several patterns such as the regular, surplus, discharge, holiday, and weekend patterns. They proposed two models to predict future usage with lower RMSE. One is the hybrid Fourier model of the Trend Switching model, which predicts usage patterns starting from the regular pattern and switches to surplus or discharge later. Another model, named the Trend Shifting model, shifts the existing pattern to match observed data. Haque et al. [14] compared four different Poisson-related models and stated that the model of Poisson with propensity and threshold-specific constant is the best. The following non-homogeneous Poisson Model proposed in [15] surpasses the other methods, such as the multivariate spatio-temporal model [16] curves similarity and historical average. Moreover, the queuing model has a non-homogeneous arrival rate and time-varying service time distribution. Hamidreza et al. [17] presented a queuing model where the arrivals follow a non-homogeneous Poisson process and predicted parking usage probabilistically from microscopic perspectives of each parking and macroscopic perspectives of aggregated parking information.
@@ II-B Machine Learning Prediction Model for General Parking Usage
Although these statistical models successfully predicted with reasonable accuracy, the models depend on sequential data such as parking availability and arrival and departure rates. Recent researchers employ machine learning techniques to integrate all related features such as weather conditions, daily patterns, and parking location and predict future usages with better accuracy [18, 19, 10]. For the research on citywide parking, Zhao and Zhang [18] developed six types of predicting models (FM1, FM2, linear regression, support vector machine (SVM), backpropagation neural network, and autoregressive integrated moving average). SVM and its ensemble model achieved lower RMSE as it is easier to find the global optimum. Zheng et al. [19] first predicted the distribution of parking arrivals and departures pattern with the Markov birth and death process. Then, with parameter estimation, they predicted parking demand for the short term. Hao et al. [10] analyzed the parking usage pattern with data aggregation at several parking sites. They found that there are tendencies in the on-working period (Sunday night to Friday daytime) and off-working period (others), in which the usage remains lower. They also developed spatio-temporal models with attention modules on the decoder and achieved high precision results. Lastly, Low et al. [20] interestingly focused on predicting the commercial vehicle parking duration. Before putting the data into the regression models, they generated the incomplete field with generative adversarial multiple imputation networks (GAMIN), then output the predicted dwell time.
@@ II-C Spatio-Temporal Prediction Model for General Parking Usage
Most prior research on predicted truck parking utilization focuses primarily on historical parking usage data and static variables, such as speed limit and annual average daily traffic [10]. However, it overlooks the inclusion of other salient predictors. As asserted by Boris and Brewster [1], the pivotal determinants guiding truck drivers’ decision-making process are parking location and the quality of amenities available. Consequently, integrating topological structures into the predictive model is justified, given its potential to account for spatial variations in the influential factors.
Graph Neural Networks (GNN) [21] are well known for understanding and representing topological structures. GNN has been widely used in the transportation field for its ability to capture spatial dependencies. A road network is naturally usable as a non-Euclidean graph structure [22]. Yang et al. [23] first leveraged Graph-Convolutional Neural Networks (GCN) on urban parking to extract the spatial dependencies of traffic flow in large-scale networks and achieved a low Mean Absolute Error (MAE) in Pittsburgh downtown. They also incorporated Recurrent Neural Networks (RNN) and Long Short Term Memory (LSTM) to capture temporal dependencies with additional related features such as weather data and traffic speed to achieve the parking usage prediction considering spatio-temporal dependencies. Zhang et al. [24] proposed a hierarchical graph convolution structure. They addressed the scarcity of real-time parking availability information with a semi-supervised approach and spatio-temporal dependencies among parking lots. Despite their success in urban parking with spatio-temporal dependencies, they have yet to conduct modeling and experiments on truck parking usage prediction due to the lack of sufficient data.
Compared to the approaches applied to citywide car parking, truck parking is more relative to their route decisions, and critical characteristics such as capacity limitations and diurnal patterns directly affect the predicted results [9]. According to Sadek et al. [9], the fixed capacity of truck parking sites represents a conductive factor in predicting truck parking usage. Besides, most truck drivers routinely work and find their rest except for unexpected situations such as bad weather or traffic congestion: the same drivers occupy most parking lots daily.
This study aims to predict future parking usage patterns considering the highway network’s topological structures and spatio-temporal dependencies of the network. Specifically, we develop a graph convolution model to capture these complex characteristics, enabled by extensive data on highway networks and parking site characteristics, such as the number of amenities and the distance from the nearest city.
@@ III-A Problem Formulation
To achieve revolving around the prediction of future truck parking usage, we employ a graph-based framework, denoted as $G=(V,E)$, where individual parking sites are represented as nodes $v\in V$, and the connections between these sites are represented as edges $e\in E$. Within this framework, we denote node embedding as $\eta$, essentially numerical representations of each site’s characteristics of features.
Our primary challenge is to predict the occupancy rates of truck parking sites over the next $T$ future time intervals. To predict the accurate future occupancy rates, we incorporate information from the $K$ preceding consecutive time intervals from the number of sites $i$. This approach enables us to capture temporal patterns in truck parking usage.
@@
$\eta_{t-K},\ldots,\eta_{t}\rightarrow\hat{\eta}_{t+1},\ldots,\hat{\eta}_{t+T}$ (1)
In this equation, the left-hand side represents the historical node embeddings for a specific parking site up to time $t$. These embeddings encode the site’s historical characteristics. Our model then processes this information to generate predictions for the occupancy rates in the subsequent time intervals, shown on the right-hand side. These predicted embeddings represent the future occupancy rates.
@@
$[X_{t-K},\ldots,X_{t};E;G]\rightarrow[\hat{X}_{t+1},\ldots,\hat{X}_{t+T}]$ (2)
To solve the task effectively, this research focuses on identifying spatial dependencies within the truck parking domain, especially regional associations between truck parking sites. Based on the preliminary analysis [12] and insights from previous studies regarding the efficacy of edge reduction techniques [25, 19], we assume that the regional relationship within the parking site network can be leveraged to reduce graph complexity. We hypothesize that a regionally decomposed graph structure, as opposed to one connected graph, offers advantages in terms of both predictive performance and computational expansiveness.
In this paper, we define smaller partitions in the target truck parking sites as an independent region and explore the spatial dependencies of each region’s site locations. Site-specific parking behavior deliberates region selection to be grounded in practicality and the unique characteristics exhibited by states [26]. For example, freight plans vary by state, so they recommend different values and strategies for developing, funding, and maintaining shared facilities. From a data aggregation perspective, it is also beneficial to manipulate data by state. Hence, our regional selection prioritizes interpretability and practical outcomes.
Figure 2 illustrates the diversity of connection approaches. We transform the connected graph into a sparser representation by incorporating both random and regional connections. Formally, we define regional graphs as subgraphs of the truck parking site graphs, denoted as $M=(V^{\prime},E^{\prime})$ segmented based on regional areas. Each regional graph $M^{r}\subseteq M$ consists of a subset of vertices $V^{r}\subseteq V$ and edges $E^{r}\subseteq E$. The adjacency matrix of the overall graph $G$ is represented as $A\in\mathbf{R}^{n\times n}$ and can be defined as follows, where the vertices of $G$ are labeled as $v_{1},\ldots,v_{n}$ and the entries are binary (${0,1}$).
@@
Similarly, the adjacency matrix of regional subgraph $M^{r}\subseteq M$ denoted as $A^{r}\subseteq A$, is selectively defined for each state in MAASTO as follows.
$e_{r_{i},r_{j}}=v_{r_{i}}v_{r_{j}}\in E^{r}\iff A^{r}[i,j]=1\$ (4)
This segregation of the graph into regional subgraphs reduces the complexity of the network in accordance with the node degrees, denoted as $d_{i}$:
$d_{r_{i}}=\sum_{j\in V_{r}}A_{r}[i,j]\subseteq\sum_{j\in V}A[i,j]$ (5)
Therefore, our framework encompasses the exploration of graph connectivity strategies, which can be categorized into three distinct approaches: (i) one connected graph, (ii) subgraphs characterized by random connections, and (iii) subgraphs constructed by the deliberate selection of nodes from regional clusters within states in the MAASTO. Consequently, our prediction task is now defined as follows.
$\begin{split}[X_{t-K},\ldots,X_{t};E;G]&=\sigma\bigcup^{r}[X_{r_{t-K+1}},\ldots,X_{r_{t}};E^{r};M^{r}]\\ &\rightarrow\bigcup^{r}[\hat{X}_{r_{t+1}},\ldots,\hat{X}_{r_{t+T}}]\end{split}$ (6)

### v2

@@ Abstract
Truck parking on freight corridors faces the major challenge of insufficient parking spaces. This is exacerbated by the Hour-of-Service (HOS) regulations, which often result in unauthorized parking practices, causing safety concerns. It has been shown that providing accurate parking usage prediction can be a cost-effective solution to reduce unsafe parking practices. In light of this, existing studies have developed various methods to predict the usage of a truck parking site and have demonstrated satisfactory accuracy. However, these studies focus on a single parking site, and few approaches have been proposed to predict the usage of multiple truck parking sites considering spatio-temporal dependencies, due to the lack of data. This paper aims to fill this gap and presents the Regional Temporal Graph Neural Network (RegT-GCN) to predict parking usage across the entire state to provide more comprehensive truck parking information. The framework leverages the topological structures of truck parking site locations and historical parking data to predict the occupancy rate considering spatio-temporal dependencies across a state. To achieve this, we introduce a Regional Decomposition approach, which effectively captures the geographical characteristics of the truck parking locations and their spatial correlations. Evaluation results demonstrate that the proposed model outperforms other baseline models, improving performance by more than $20\%$.
@@ I Introduction
The Federal Motor Carrier Safety Administration (FMCSA) has implemented regulations to reduce fatigue-related safety hazards for truck drivers by restricting the consecutive and total daily driving hours permitted for truck operators. Despite these efforts, truck drivers still spend considerable time searching for safe parking locations beyond their limited legal working hours [1]. This phenomenon highlights the persistent issue of inadequate truck parking facilities, contributing to unauthorized parking behaviors and breaches of parking regulations [2].
@@
The literature identifies increasing truck drivers’ accessibility to real-time parking usage information [3] and drivers’ perceptual interpretations [4, 5] as major factors influencing such parking challenges. According to the survey studies in the region [6], members of the Mid-America Association of State Transportation Officials (MAASTO) from eight states, Indiana, Iowa, Kansas, Kentucky, Michigan, Minnesota, Ohio, and Wisconsin, have collaborated to build a real-time multistate Truck Parking Information Management System (TPIMS) to provide more practical truck parking information to truck drivers [7, 8]. TPIMS, fully operational since January 4, 2019, provides real-time parking usage information to truck drivers through dynamic message signs, smartphone applications, traveler information websites (e.g. 511 traveler information) and other forms.
@@
In addition to real-time parking site availability information, future usage prediction can better support truck drivers in their decision-making process and route planning. Current prediction models exhibit commendable accuracy in predicting future usage for the subsequent hour [9, 10]. However, it is crucial to recognize that the domain of interest encompasses multiple truck parking sites situated along the highway corridor. Boris and Brewster [1] reported that the most critical aspects of the target parking sites are the proximity to the route/destination followed by the availability of amenities. Therefore, the primary objective of our prediction model is to capitalize on the wealth of data available from various truck parking sites and leverage their topological structures.
@@
Since 2019, the research team has developed the TPIMS data archiving system (TPIMS-DAS), designed to store effectively and track truck parking information [11]. Leveraging the substantial benefits derived from an abundance of comprehensive and detailed information on truck parking sites, users can extract and aggregate truck parking data from eight states in MAASTO using TPIMS-DAS. It also provides a sufficient amount of truck parking data to analyze truck parking usage patterns [12] and train reliable models. In this research, we aim to develop a data-driven, predictive model for future parking usage, grounded in spatio-temporal dependencies manifested by geographical locations of parking sites on highway corridors and their historical usage patterns.
@@
To this end, we propose a Regional Temporal Graph Neural Network (RegT-GCN), comprising a Graph Neural Network with temporal capability and a novel decomposition module, Regional Decomposition, to capture regional/spatial relationships (Figure 1). In this study, a region is defined as a state to incorporate their unique characteristics (e.g., state laws). The significant historical parking data allows the model to capture temporal dependencies of multi-state-wide truck parking usage to generate more accurate prediction results. We conduct several experiments to evaluate the model against various state-of-the-art models and for the importance of Regional Decomposition for Graph Neural Networks.
Our contributions are summarized as follows.
  • A spatio-temporal model to predict the usage of multiple truck parking sites.
@@
  • Extensive quantitative evaluations, which demonstrate that our RegT-GCN significantly outperforms state-of-the-art baseline models.
@@ II Literature Review
Parking prediction methods can be classified into two main categories: urban parking prediction, which is related to parking prediction within cities, and truck parking prediction, which deals with predicting parking usage for trucks. Typically, urban parking prediction methods offer a shorter prediction horizon, ranging from 5 to 15 minutes ahead. In contrast, truck parking prediction provides predictions for up to 1 hour ahead [10]. Regardless of the distinctions in time frames between these parking prediction tasks, it should be noted that both tasks often share fundamental concepts and are devised with a broader perspective that encompasses diverse parking scenarios.
@@ II-A Statistical Prediction Model for Truck Parking Usage
Bayraktar et al. [13] developed an intelligent parking-management system for commercial vehicles in Florida, analyzed data, and created a prediction model for rest area occupancy with a Kalman filter. As the Kalman filter prediction-correction mechanism can be utilized with historical data and adapted to the online status, its model achieved better performance. Sadek et al. [9] identified several patterns such as the regular, surplus, discharge, holiday, and weekend patterns. They proposed two models to predict future usage with lower RMSE. One is the hybrid Fourier model of the Trend Switching model, which predicts usage patterns starting from the regular pattern and switches to surplus or discharge later. Another model, named the Trend Shifting model, shifts the existing pattern to match the observed data. Haque et al. [14] compared four different Poisson-related models and stated that the Poisson model with propensity and threshold-specific constant is the best. The following non-homogeneous Poisson model proposed in [15] surpasses the other methods, such as the multivariate spatio-temporal model [16] curve similarity and historical average. Moreover, the queuing model has a non-homogeneous arrival rate and time-varying service time distribution. Hamidreza et al. [17] presented a queuing model in which arrivals follow a non-homogeneous Poisson process and predicted parking usage probabilistically from microscopic perspectives of each parking and macroscopic perspectives of aggregated parking information.
@@ II-B Machine Learning Prediction Model for General Parking Usage
Although these statistical models successfully predicted with reasonable accuracy, the models depend on sequential data such as parking availability and arrival and departure rates. Recent researchers have used machine learning techniques to integrate all related features such as weather conditions, daily patterns, and parking location and predict future usages with better precision [18, 19, 10]. For the research on citywide parking, Zhao and Zhang [18] developed six types of predicting models (FM1, FM2, linear regression, support vector machine (SVM), backpropagation neural network, and autoregressive integrated moving average). SVM and its ensemble model achieved lower RMSE as it is easier to find the global optimum. Zheng et al. [19] first predicted the distribution of the parking arrival and departure patterns with the Markov birth-and-death process. Then, with parameter estimation, they predicted the parking demand for the short term. Hao et al. [10] analyzed the parking usage pattern with data aggregation at several parking sites. They found that there are tendencies in the on-working period (Sunday night to Friday daytime) and off-working period (others), in which the usage remains lower. They also developed spatio-temporal models with attention modules on the decoder and achieved high-precision results. Lastly, Low et al. [20] interestingly focused on predicting the duration of commercial vehicle parking. Before putting the data into the regression models, they generated the incomplete field with generative adversarial multiple imputation networks (GAMIN), then output the predicted dwell time.
@@ II-C Spatio-Temporal Prediction Model for General Parking Usage
Most prior research on predicted truck parking utilization focuses primarily on historical parking usage data and static variables, such as speed limit and annual average daily traffic [10]. However, it overlooks the inclusion of other significant predictors. According to Boris and Brewster [1], the key determinants guiding truck drivers’ decision-making process are parking location and the quality of the available amenities. Consequently, integrating topological structures into the predictive model is justified, given its potential to account for spatial variations in the influential factors.
Graph Neural Networks (GNN) [21] are well known for understanding and representing the topological structures of the input. GNN has been widely used in the transportation domain for its ability to capture spatial dependencies. A road network is naturally functional as a non-Euclidean graph structure [22]. Yang et al. [23] first leveraged Graph-Convolutional Neural Networks (GCN) on urban parking to extract the spatial dependencies of traffic flow in large-scale networks and achieved a low Mean Absolute Error (MAE) in Pittsburgh downtown. They also incorporated Recurrent Neural Networks (RNN) and Long Short Term Memory (LSTM) to capture temporal dependencies with additional related features such as weather data and traffic speed to achieve the parking usage prediction with regard to spatio-temporal dependencies. Zhang et al. [24] proposed a hierarchical graph convolution structure. They addressed the scarcity of real-time parking availability information with a semi-supervised approach and spatio-temporal dependencies among parking lots. Despite their success in urban parking with spatio-temporal dependencies, they have yet to conduct modeling and experiments on truck parking usage prediction due to the lack of sufficient data.
Compared to approaches applied to urban car parking, truck parking is more relative to truck drivers’ route decisions, and critical characteristics such as the existence of amenities, capacity limitations, and diurnal patterns directly affect the predicted results [9]. The parking selection strategies of drivers exhibit distinction, with truck drivers typically opting for parking site along highways while car drivers predominantly seek parking proximate to their urban destinations. Furthermore, most truck drivers routinely work and find their resting place during night or weekends, different from the temporal patterns observed in urban parking scenarios, where drivers look for parking during the day [25].
This study aims to predict future parking usage patterns considering the highway network’s topological structures and spatio-temporal dependencies of the network. Specifically, we develop a graph convolution model to capture these complex characteristics, enabled by extensive data on highway networks and parking characteristics, such as the number of amenities and the distance from the nearest city.
@@ III-A Problem Formulation
We describe the problem formulation for the task of predicting future truck parking usage, as a novel model to represent spatio-temporal dependencies on the parking. Starting from the general notation, a truck parking graph is denoted as $G=(V,E)$, where individual parking sites are represented as nodes $v\in V$, and the connections between these sites are represented as edges $e\in E$. Within this framework, we denote node-embedding as $\eta$, representing each site’s characteristics of features numerically.
Our primary challenge is to predict the occupancy rates of truck parking sites in the next $T$ future time intervals. To predict accurate future occupancy rates $o$ at the site $i$, we incorporate information from the $K$ preceding consecutive time intervals. This approach enables us to capture temporal patterns in the use of truck parking.
@@
$\eta_{i_{t-K}},\ldots,\eta_{i_{t}}\rightarrow\hat{\eta}_{i_{t+1}},\ldots,\hat{\eta}_{i_{t+T}}$ (1)
In this equation, the left-hand side represents the historical node embeddings for a specific parking site up to time $t$. These embeddings encode the site’s historical characteristics. Our model then processes this information to generate predictions for the occupancy rates in the subsequent time intervals, shown on the right-hand side. These predicted embeddings represent future occupancy rates.
@@
$[V_{t-K},\ldots,V_{t};E;G]\rightarrow[\hat{V}_{t+1},\ldots,\hat{V}_{t+T}]$ (2)
In order to capture the spatio-temporal dependencies, this research first engages in understanding regional relationships across truck parking sites. Based on our preliminary analysis [12] and insights from previous studies on the efficacy of edge reduction techniques [19], [26], we assume that the regional relationship within the truck parking network can be leveraged to reduce graph complexity. We hypothesize that a regionally decomposed graph structure, as opposed to one connected graph, offers advantages in terms of both predictive performance and computational efficiency.”
In this paper, we define small partitions in the truck parking sites as an independent region and explore the spatial dependencies among truck parking sites: state. Site-specific parking behavior deliberates region selection to be grounded in practicality and the unique characteristics exhibited by states [26]. For example, freight plans vary by state, so they recommend different values and strategies for developing, funding, and maintaining shared facilities. It is also beneficial to consider the state as a region from a data aggregation perspective. Hence, our regional selection prioritizes interpretability and practical outcomes.
Formally, we define regional graphs as subgraphs of truck parking graphs, denoted as $M=(V^{\prime},E^{\prime})$ segmented according to regional areas, where $M\neq G,M\subseteq G$. Each regional graph $M^{r}\subseteq M$ consists of a subset of vertices and edges $M^{r}=\{(V^{r},E^{r})|\forall r\in\text{region}\}$. The adjacency matrix of the overall graph $G$ is represented as $A\in\mathbf{R}^{N\times N}$ and can be defined as follows, where the vertices of $G$ are labeled $v_{1},\ldots,v_{n}$ and the entries are binary (${0,1}$).
@@
Similarly, the adjacency matrix of the regional subgraph $M^{r}\subseteq M$ denoted as $A^{r}\subseteq A$, is selectively defined for each state in MAASTO as follows.
$e^{r}_{i,j}=v^{r}_{i}v^{r}_{j}\in E^{r}\iff A^{r}[i,j]=1$ (4)
This segregation of the graph into regional subgraphs reduces the complexity of the network according to the degrees of the nodes, denoted as $d^{i}_{r}$ for site $i$:
$d^{r}_{i}=\sum_{j\in V^{r}}A^{r}[i,j]\subseteq\sum_{j\in V}A[i,j]$ (5)
Therefore, our framework incorporates the invention of graph decomposition strategies with the construction of subgraphs by deliberately selecting nodes from regional sets within states in the MAASTO. Our prediction task is now defined as follows.
$\begin{split}[V_{t-K},\ldots,V_{t};E;G]&=\sigma\bigcup^{r}[V^{r}_{t-K+1},\ldots,V^{r}_{t};E^{r};M^{r}]\\ &\rightarrow\bigcup^{r}[\hat{V}^{r}_{t+1},\ldots,\hat{V}^{r}_{t+T}]\end{split}$ (6)